OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (4, -7) and a directrix of y = -15. Put the equation in standard form

OpenStudy (anonymous):

a. f(x) = 1/16 x2 - 8x + 11 b. f(x) = 1/16 x2 - 8x - 10 d. f(x) = 1/16 x2 - 1/2 x - 10

OpenStudy (anonymous):

C was wrong

OpenStudy (campbell_st):

well you need to find the vertex. but before that you need to find the focal length the distance between the directrix and focus is equal to 2 times the focal length. so what do you think the focal length would be..?

OpenStudy (anonymous):

I'm totally lost! The practice problems don't make sense at all:(

OpenStudy (anonymous):

And in the lesson it didn't say anything about needing the vertex or focal length I don't think

OpenStudy (campbell_st):

ok...  the vertex in the line x = 4 and located half way between y = -7 and y = -15 the total distance between y = -7( the focus) and y = -15( the directrix) is a length which is double the focal length